what's Bills current age and Bills nephews current age if he is 4 times older than his nephew and in 5 years the sum of their ages will be 50?
Answer by ankor@dixie-net.com(22740)

what's Bills current age and Bills nephews current age if he is 4 times older than his nephew and in 5 years the sum of their ages will be 50?
:
let n = nephews current age
Bill is 4 times older,therefore:
4n = Bills current age
:
In five years
(n+5) + (4n+5) = 50
n + 4n + 5 + 5 = 50
5n + 10 = 50
5n = 50 - 10
5n = 40
n = 40/5
n = 8 yrs is Nephews age
then obviously
4(8) = 32 yrs is Bills age
:
:
Confirm this in the statement
"in 5 years the sum of their ages will be 50?"
(8+5) + (32+5) =
13 + 37 = 50